Question 1 The adjusted trial balance for Marios Pizza Palace at 30 June is presented below: Marios Pizza Palace Adjusted tDelivery income 430 32 500 Food purchases 55,310 510 Purchase returns and allowances 418 13,652 520 Packaging costs DiscountCash No. 100 Post Post Date Details Debit Date Details Credit ref. ref. 32,957 June 30 June 1 Balance Payments 12,699 CPJ12 RInventory - packaging No. 140 Post Post Credit Date Details Debit Date Details ref. ref. x...........x X....... 921 June June15578 Accounts payable No. 200 Post Post Date Details Debit Date Details Credit ref. ref. 330 June 1 June 15 Purchase returnsNEXT P Bank loan No. 280 Post Post ref. Details Debit Details Date Date Credit ref. 1,030 June 1 Cash 19,836 June 30 BalanceSales income No. 400 Post Post Credit Date Details Debit Date Details ref. ref. x........ 163,473 30 + 183561 Balance June JuPurchase returns and allowances No. 510 Post Post ref. Date Details Debit Date Details Credit ref. ........... 88 30 : 418 JuBank charges No. 610 Post ref. Post Date Details Debit Date Details Credit ref. 30 : 861 799 June June 1 Balance 62 30 Cash C

Fill in the blanks
Answer #1

Please go though above templates and trial balance to understand the answer provided below:-

1. In cash ledger shown in first screenshot of template above will be filled as follows

As trial balance as on 30th june is available, cash balance shown in trial balance is 40290 and same is filled as balance c/d on credit side in cash ledger. We have to fill the b/f balance on 1 july, it will be same as of closing balance on 30th june i.e. 40290.

In the same way, blank provided in accounts receivable account with 668.

2. Inventory shown in trial balance will be opening stock and closing stock value will be available as a note to trial balance.

Balance on 30th june for inventory-food will be 1855 as provided in notes and same will be shown in amount b/f on 1st july. Another blank column on credit side of inventory account will be filled with balancing figure amount i.e. 1205.

In the same way inventory packaging ledger can be filled.

3. Prepaid advertising- closing balance is shown in trial balance as 288 and the blank below shows the total so it will be filled with 288+288=576

Similarly, prepaid insurance blank will be filled.

4. Blanks provided on both sides in accounts payable ledger are total i.e 330+2842+13781=2262+14400+291=16953

5. Owner's capital closing balance is 59676 as provided in trial balance and other blank on credit side will the profit for the period i.e.20507(calculated in pt. no. 6. below and blank on debit side will be calculated as balancing figure i.e.59676+20507-41712=38471.

Blank below this blank is of total i.e. 59676+20507=80183.

6. In profit & loss summary, income will be 183561-49+537+32=184081

Purchase return/ inventory = 418+1855+3893=6166

Expenses will be total of items shown in trial balance from food purchases to wage expenses i.e. 55310+13652+274+77+1278+3798+861+6958+3028+2766+1005+5224+426+26776+716+348+3313+3790+40140=169740

Debit side blank column will be filled with net profit for the period= 184081+6166-169740=20507

7. In all the remaining ledgers, the blank given under detail columns are to be filled with word "Profit & loss account".